# The effect of control barrier functions on energy transfers in controlled physical systems

### v1

@@ Abstract
Using a port-Hamiltonian formalism, we show the qualitative and quantitative effect of safety-critical control implemented with control barrier functions (CBFs) on the power balance of controlled physical systems. The presented results will provide novel tools to design CBFs inducing desired energetic behaviors of the closed-loop system, including nontrivial damping injection effects and non-passive control actions, effectively injecting energy in the system in a controlled manner. Simulations validate the stated results.
Index Terms:
Control barrier functions, port-Hamiltonian systems, energy-based control, nonlinear control
@@ I Introduction
Control objectives in engineering applications are becoming increasingly complex in terms of their semantic and mathematical description. An instructive example to think about is safe collaboration between humans and robots [Robla-Gomez2017WorkingEnvironments, Hamad2023AInteraction, Siciliano2016SpringerRobotics], where control goals are often represented by multi-objective functions representing the specific task to be executed. These objectives (which are sometimes unknown and implicitly designed through e.g. control by human demonstrations) are non-trivially intertwined with safety specifications, aimed at representing critical hazards whose prevention needs to be certified before any tentative of real-world application. These safety specifications are often conflicting with the main task to be executed, and the semantic distinction between the “main task”, represented by objective functions, and the “safety specifications”, represented by constraints on the state of the system in the optimization problem, is ubiquitous in the control-theoretic literature [Ames2019ControlApplications, Wabersich2023Data-DrivenSystems, Ferraguti2022SafetyApproach]. This view of safety-critical control applications has led to the concept of safety filters [Wabersich2023Data-DrivenSystems], which are computational units with the goal of transforming a nominal control input (designed for the main task, disregarding the safety requirements) into a new control input which i) minimally modifies the nominal controller and ii) ensures the safety requirements. One of the most popular techniques for implementing these safety filters is represented by control barrier functions (CBF) [Ames2019ControlApplications, Ames2017ControlSystems], an algorithm which aims to modify a nominal control input to achieve forward invariance of a safe set.
@@
In this paper, we perform an in-depth analysis of the effect of control barrier functions (CBFs) on energy transfers in physical systems, inducing novel control design tools aimed at regulating the energy transfers along the closed-loop system. The idea that an energetic description captures the behavioral property of a system is present in some branches of control theory, especially in the so-called energy-based or energy-aware control literature [Duindam2009ModelingSystems, Stramigioli2015Energy-AwareRobotics, Califano2022OnSystems, Ortega2001PuttingControl]: if we are able to control the energy transfers undergoing the controlled system, then we can achieve desired behaviors, going beyond mere stabilization purposes. This aspect is particularly valid in collaborative robotic applications that take place in unstructured environments, possibly in the presence of humans, since energy transfers between the robot and its surroundings encode information on performance, safety and energy efficiency at the same time [Tadele2014CombiningRobots, Laffranchi2009SafeControl, Zacharaki2020SafetySurvey, Califano2022OnSystems, Tadele2014ThePublications]. This narrative of “putting energy back in control” [Ortega2001PuttingControl] has been the driving motivation in the development of control schemes in the port-Hamiltonian formalism [Duindam2009ModelingSystems].
Port-Hamiltonian (pH) systems are control-affine systems that serve best to represent autonomous and controlled physical systems, since their structural properties directly relate to their underlying energetic structures. PH systems basically encompass all physical systems of interest, finite- or infinite-dimensional, and from any physical domain [Duindam2009ModelingSystems, Rashad2020TwentyReview]. As a consequence, the use of this formulation does not introduce any restriction to the class of physical systems that can be modeled, while at the same time it introduces technical and intuitive advantages when studying the effects of CBF algorithms on the energy balance of the system. Among the possible energy-aware control strategies that can be developed with the tools presented in this work, we will present novel damping injection schemes, in which the CBF algorithm induces passive control actions that extract energy from the controlled system. Furthermore, we introduce new schemes which effectively inject energy in the controlled system, going beyond passive designs and giving new perspectives to achieve desired closed-loop behaviors.
@@
In Sec. II the background on pH systems and CBFs is reviewed. In Sec. III the main contribution of this letter is given and applied to mechanical systems in Sec. IV. In Sec. V we show numerical simulations and Sec. VI contains conclusions and future work.
@@ II Background
We refer to [vanderSchaftL2, Duindam2009ModelingSystems] regarding pH systems and to [Ames2019ControlApplications, Xu2015RobustnessControl, Ames2017ControlSystems, Singletary2021Safety-CriticalSystems] regarding CBF for references that fully cover the presented background. The specialisation of CBFs to pH systems is a contribution on its own present in this section.
@@ II-A Port-Hamiltonian systems
where $x\in\mathcal{D}\subseteq\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the input, $g(x)$ is the input matrix, $J(x)=-J(x)^{\top}$ and $R(x)=R(x)^{\top}\geq 0$ are respectively skew-symmetric and positive semi-definite symmetric matrices representing the power-preserving and the dissipative components of the system. The non-negative function $H:\mathcal{D}\to\mathbb{R}^{+}$ is called the Hamiltonian and maps the state into the total physical energy of the system. $\partial_{x}H(x)\in\mathbb{R}^{n}$ denotes the gradient of $H$, represented as a column, and $\partial_{x}^{\top}H(x)$ denotes its transposed. When not explicitly stated, all described variables are assumed to have a degree of continuity such that the right-hand side of (1) is locally Lipschitz, to guarantee the existence and uniqueness of the solutions.
We now present geometric properties of port-Hamiltonian systems that are relevant in this paper. The use of coordinate-free notation will induce intuitive and technical advantages in the interpretability of the expressions and in the calculations.
@@ II-A3 Passivity
The latter power balance is a statement of energy conservation for the physical system (1), where it should be appreciated how the pH structure displays all the possible ways in which physical energy can flow along the system: the skew-symmetric structure $J$ represents pure routing of energy, the dissipative structure $R$ dissipated energy, and the duality product $y^{\top}u$ represents the instantaneous power injected by the input. Passivity condition (2) states that the variation of energy in the system is bounded by the injected power.
@@ II-A4 Control
From a system-theoretic perspective passivity implies stability under weak conditions: $H(x)$ is a valid Lyapunov candidate since $u=0$ implies $\dot{H}\leq 0$. Since (1) is time-invariant, the autonomous system converges to the largest invariant subset of $\{x\in\mathcal{D}|[H,H]_{R}=0\}$ which generally depends on the dissipation matrix $R(x)$ and the Hamiltonian $H(x)$. Many pH inspired control schemes have been proposed, with the rationale of obtaining a closed-loop system in the form (1) with Hamiltonian and system matrices corresponding to desired behaviors.
@@ II-B Control barrier functions
The safe set $\mathcal{C}$ is built as the superlevel set of a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$, i.e., $\mathcal{C}=\{x\in\mathcal{D}:h(x)\geq 0\}.$ The function $h(x)$ is then defined as a CBF on $\mathcal{D}$ if $\partial_{x}h(x)\neq 0,\forall x\in\partial\mathcal{C}$ and
@@
for all $x\in\mathcal{D}$ and some extended class $\mathcal{K}$ function [^1] A function $\alpha:(-b,a)\to(-\infty,\infty)$ with $a,b>0$, which is continuous, strictly increasing, and $\alpha(0)=0$. $\alpha$. Here we used $\dot{h}=\partial_{x}^{\top}h(x)\dot{x}$ to denote the time derivative of $h$ along the solution of (1)), which results in:
@@
The link between the existence of a CBF and the forward invariance of the related safe set is established by the following key result.
Theorem 1 ([Ames2017ControlSystems], here applied to pH systems (1)).
Let $h(x)$ be a CBF on $\mathcal{D}$ for (1). Any locally Lipschitz controller $u=k(x)$ such that $\{h,H\}_{J}-[h,H]_{R}+\partial_{x}^{\top}h(x)g(x)k(x)\geq-\alpha(h(x))$ provides forward invariance of the safe set $\mathcal{C}$. Additionally the set $\mathcal{C}$ is asymptotically stable on $\mathcal{D}$.
@@
The transformation of the desired control input $u_{\textrm{nom}}(x)$ in $u^{*}(x)$ by solving (6) is denoted as safety-critical control. A last result that will be crucially used in this work is the following lemma.
Theorem 2 ([Xu2015RobustnessControl, Singletary2021Safety-CriticalSystems], here applied to pH systems (1)).
Let $h(x)$ be a CBF on $\mathcal{D}$ for (1) and assume $\mathcal{U}=\mathbb{R}^{m}$. Define $\Psi(x):=\dot{h}(x,u_{\textup{nom}}(x))+\alpha(h(x))$. Omitting functional dependencies for readability, a closed-form solution for (6) is given by $u^{*}(x)=u_{\textup{nom}}(x)+u_{\textup{safe}}(x)$, where
$u_{\textup{safe}}(x)=\begin{cases}\begin{aligned} &-\frac{g^{\top}\partial_{x}h}{[h,h]_{gg^{\top}}}\Psi&\textup{if}\,\Psi<0\\ &\qquad\qquad 0&\textup{if}\,\Psi\geq 0\end{aligned}\end{cases}.$ (7)
For the remainder of this paper we will assume $u_{\textrm{nom}}=0$. This choice is consistente with the way controllers are designed in the pH framework. In fact, as reported in Sec. II, the form (1) can already represent controlled physical systems performing some nominal task with a desired energetic behavior.
@@ III Putting energy back in control through CBFs
In this section, we discuss how CBF-induced safety-critical control influences the power balance of the controlled physical system. We start presenting a closed-form solution for the power injected by the safety-critical controller for controlled pH systems.
@@
Let the assumptions of Theorem 2 hold. The power balance for the pH system (1), controlled with input $u_{\textrm{nom}}(x)$ undergoing safety-critical control (6) induced by a CBF $h(x)$, is
$\dot{H}=-[H,H]_{R}-\mathbbm{1}_{\Psi<0}\frac{[H,h]_{gg^{\top}}}{[h,h]_{gg^{\top}}}\Psi,$ (8)
where $\Psi=\{h,H\}_{J}-[h,H]_{R}+\alpha(h(x))$ and the indicator function $\mathbbm{1}_{\Psi<0}$ returns $1$ if $\Psi<0$ ($u_{\textrm{safe}}\neq 0$) and $1$ otherwise ($u_{\textrm{safe}}=0$).
@@
The results follows from an explicit calculation of the power balance (2) using $u=u_{\textup{safe}}(x)$ given by (7). ∎
@@
$P_{h,\alpha}=-\mathbbm{1}_{\Psi<0}\frac{[H,h]_{gg^{\top}}}{[h,h]_{gg^{\top}}}\Psi$ (9)
where the notation stresses the fact that this term depends both on the CBF $h$ and on the function $\alpha$, a parameter of the CBF algorithm whose role will be understood in energetic terms in the sequel.
III-A Preserving stability: CBF-based damping injection
As reported in Sec. II-A4, if the Hamiltonian $H$ qualifies as a Lyapunov function, the minimum of $H$ is a stable equilibrium (the time derivative of $H$ is given by (8) with $P_{h,\alpha}=0$). Natural questions that can be answered are: under what conditions does the safety-critical control preserve the stability of the underlying controlled system? If stability is preserved, what is the effect of safety-critical control on the rate of convergence to the equilibrium? As (8) holds true, in case $H$ would still qualify as a Lyapunov function after the safety-critical control, the closed-loop system would converge to the largest invariant subset in $\{x\in\mathcal{D}|P_{h,\alpha}-[H,H]_{R}=0\}$. The following result gives a useful sufficient condition for preservation of local stability.
@@ III-A Preserving stability: CBF-based damping injection
Suppose $H$ is a Lyapunov function for the pH system (1). Suppose that a safety-critical control algorithm induced by a CBF $h(x)$ acts on the system under the assumptions of Theorem 2. The stability of the original system is preserved if $\mathbbm{1}_{\Psi<0}[H,h]_{gg^{\top}}\leq 0$ in a neighborhood of the equilibrium.
@@
The result follows by working out analitically the sufficient condition under which $P_{h,\alpha}$, and as such the whole right-hand side of (8), is non-positive. ∎
@@
The condition $P_{h,\alpha}\leq 0$ forces the safety-critical controller to “act as a damper”, possibly helping dissipative dynamics encoded in (1) in the stabilization task, normally referred to as damping injection in the pH framework. The closed-form expression (9) can be used to design and interpret the non-trivial damping effects that specific CBFs have on the controlled system.
III-B Beyond damping injection
In a safety-critical framework, we find a strong conceptual distinction between the “main task” to be executed and the “safety constraints” which have to be guaranteed in any circumstance. Although this narrative is useful in many applications and is widespread in the control-theoretic literature, the main controller and the CBF-induced term must not necessarily have antagonistic roles, as seen e.g., in previous section where the safety-critical controller can act as a smart damping injection mechanism. More in general, the power term induced by CBFs (9) induces a constructive way to “put energy back in control”, in the sense envisioned in energy-based methods (see e.g., [Ortega2001PuttingControl] and [Stramigioli2015Energy-AwareRobotics]), also going beyond purely passive designs and possibly injecting positive power in the system if the task requires it. To show the potential of the approach in the context of pH systems, it is instructive to design a CBF that encodes as a safe set the fact that the system has to stay below or above a certain energy limit $c$, that is, $h(x)=\pm(c-H(x))$. The following example shows what a control designer can gain in terms of interpretability of the safety-critical control action, as well as the role of the class $\mathcal{K}$ function $\alpha$ in the algorithm.
Example 1 (Limiting total energy from above and below).

### v2

@@ Abstract
Using a port-Hamiltonian formalism, we show the effects of safety-critical control implemented with control barrier functions (CBFs) on the power balance of controlled physical systems. The presented results will provide novel tools to design CBFs inducing desired energetic behaviors of the closed-loop system, including non-trivial damping injection effects and non-passive control actions, effectively injecting energy into the system in a controlled manner. Simulations validate the presented results.
@@ I Introduction
Control objectives in engineering applications are becoming increasingly complex in terms of their semantic and mathematical description. An instructive example to think about is safe collaboration between humans and robots [Robla-Gomez2017WorkingEnvironments, Hamad2023AInteraction, Siciliano2016SpringerRobotics], where control goals are often represented by multi-objective functions representing the specific task to be executed. These objectives (often unknown or implicitly designed through e.g. control by human demonstrations) are non-trivially intertwined with safety specifications, aimed at representing critical hazards whose prevention needs to be certified before any attempt at real-world application. These safety specifications are often conflicting with the main task to be executed, and the semantic distinction between the “main task", represented by objective functions, and the “safety specifications", represented by constraints on the state of the system in the optimization problem, is ubiquitous in the control-theoretic literature [Ames2019ControlApplications, Wabersich2023Data-DrivenSystems, Ferraguti2022SafetyApproach]. This view of safety-critical control applications has led to the concept of safety filters [Wabersich2023Data-DrivenSystems], which are computational units with the goal of transforming a nominal control input (designed for the main task, disregarding the safety requirements) into a new control input which i) minimally modifies the nominal controller and ii) ensures the safety requirements. One of the most popular techniques for implementing these safety filters is represented by control barrier functions (CBF) [Ames2019ControlApplications, Ames2017ControlSystems], an algorithm which aims to modify a nominal control input to achieve forward invariance of a safe set.
@@
In this paper, we perform an analysis of the effect of CBFs on energy transfers in physical systems, and present novel control tools aimed at regulating the energy in the closed-loop system. The idea that an energetic characterization captures high-level behavioral properties of a system is present in the so-called energy-based or energy-aware control literature [Duindam2009ModelingSystems, Stramigioli2015Energy-AwareRobotics, Califano2022OnSystems, Ortega2001PuttingControl]: the ability to control the energy transfers in controlled system induces desired behaviors, beyond mere stabilization purposes. Energy-aware designs serve to achieve high level control objectives based on energy, including and extending passivity-based control (PBC) strategies [vanderSchaftL2], which are used to robustly achieve stabilization goals. Energy-aware control is particularly interesting in collaborative robotic applications that take place in unstructured environments, possibly in the presence of humans, since energy transfers between the robot and its surroundings encode information on performance, safety, and energy efficiency [Tadele2014CombiningRobots, Laffranchi2009SafeControl, Zacharaki2020SafetySurvey, Califano2022OnSystems, Tadele2014ThePublications]. Another instructive example embodying the need of an energy-aware design is periodic locomotion tasks, as purely passive designs cannot sustain them without continuous energy injection to counteract internal dissipation. The narrative of “putting energy back in control" [Ortega2001PuttingControl] has been the driving motivation in the development of control schemes in the port-Hamiltonian (pH) formalism [Duindam2009ModelingSystems, Rashad2020TwentyReview]. PH systems encodes energetic properties of autonomous and controlled physical systems, and will be conveniently used in this paper to gain intuitive and technical advantages in the energetic analysis. While energy-aware methods claim to manage energy flow, practical tools to realize this are still missing. This work provides a novel conceptual and methodological approach to energy-aware control, where CBFs are designed to actively regulate energy flows, including controlled energy injections into the closed-loop system.
We recognize other works combining the CBF methodology with energetic analysis of the closed-loop system. In [ROMDLONY201639] the authors combine pH modeling and CBFs to obtain a stable closed-loop system with safety guarantees. In [Capelli2022PassivityEnergy] CBFs are used in combination with energy tanks [Califano2022OnSystems] to passivize a potentially non passive control action. In [Notomista2019Passivity-BasedFunctions] a time varying CBF is designed to passivize a closed-loop system. [10787049] develops an optimization that makes the closed-loop system passive and safe at the same time. In [10136379] a class of CBFs preserving the passivity of a closed-loop system were characterized with technical tools similar to those used in this paper. [10886724] implements a similar idea for switching systems. Other relevant works are [Singletary2021Safety-CriticalSystems], which introduces so called energy-based CBFs to deal with kinematic constraints in robotics, and [9561981], where a CBF-inspired controller is developed in a similar context and comprising a passivity and safety analysis. In all these works the desired energetic properties of the closed-loop system were limited to passivity, considered as a necessary condition for some form of safety. In this work we will extend the use of CBFs to energy-aware designs. In particular our contributions include the introduction of CBFs along: i) damping injection schemes, in which energy is extracted from the controlled system, with novel sufficient conditions on closed-loop stability; and ii) new schemes which effectively inject energy in the controlled system, going beyond passive designs and giving new perspectives to achieve desired closed-loop energetic behaviors. This last contribution is validated through a task resembling periodic locomotion, introducing a completely new perspective on the use of CBFs.
@@
The paper is organized as follows. In Sec. II the background on pH systems and CBFs is reviewed. In Sec. III the main contribution of this letter is given and applied to mechanical systems in Sec. IV. In Sec. V we show numerical simulations and Sec. VI contains conclusions and future work.
@@ II Background
We refer to [vanderSchaftL2, Duindam2009ModelingSystems] regarding pH systems and to [Ames2019ControlApplications, Xu2015RobustnessControl, Ames2017ControlSystems, Singletary2021Safety-CriticalSystems] regarding CBF for references that fully cover the presented background.
@@ II-A Port-Hamiltonian systems
where $x\in\mathcal{D}\subseteq\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the input, $g(x)$ is the input matrix, $J(x)=-J(x)^{\top}$ and $R(x)=R(x)^{\top}\geq 0$ are respectively skew-symmetric and positive semi-definite symmetric matrices representing the power-preserving and the dissipative components of the system. The non-negative function $H:\mathcal{D}\to\mathbb{R}^{+}$ is called the Hamiltonian and maps the state into the total physical energy of the system. $\partial_{x}H(x)\in\mathbb{R}^{n}$ denotes the gradient of $H$, represented as a column, and $\partial_{x}^{\top}H(x)$ denotes its transposed. When not explicitly stated, all described variables are assumed to have a degree of continuity such that the right-hand side of (1) is locally Lipschitz, to guarantee the existence and uniqueness of the solutions. We now present geometric properties of port-Hamiltonian systems that are relevant in this paper.
@@ II-A3 Passivity
The latter power balance is a statement of energy conservation for the physical system (1), where all ways in which physical energy can flow along the system are displayed: the skew-symmetric structure $J$ represents pure routing of energy, the dissipative structure $R$ dissipated energy, and the duality product $y^{\top}u$ represents the instantaneous power injected by the input. Passivity condition (2) states that the variation of energy in the system is bounded by the injected power.
@@ II-A4 Control
From a system-theoretic perspective passivity implies stability under weak conditions: $H(x)$ is a valid Lyapunov candidate since $u=0$ implies $\dot{H}\leq 0$. Since (1) is time-invariant, the autonomous system converges to the largest invariant subset of $\{x\in\mathcal{D}|[H,H]_{R}=0\}$ which generally depends on the dissipation matrix $R(x)$ and the Hamiltonian $H(x)$. Many pH-inspired control schemes have been proposed, with the rationale of obtaining a closed-loop system in the form (1) with Hamiltonian and system matrices corresponding to desired behaviors. Importantly, IDA-PBC designs [ORTEGA2002585], indeed aim at obtaining a closed-loop port-Hamiltonian system encoding desired energetic properties.
@@ II-B Control barrier functions
The safe set $\mathcal{C}$ is built as the superlevel set of a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$, i.e., $\mathcal{C}=\{x\in\mathcal{D}:h(x)\geq 0\}.$ The function $h(x)$ is then defined as a CBF on $\mathcal{D}$ if $\partial_{x}h(x)\neq 0,\forall x\in\partial\mathcal{C}=\{x\in\mathcal{D}:h(x)=0\}$, and
@@
for all $x\in\mathcal{D}$ and some extended class $\mathcal{K}$ function [^1] A function $\alpha:(-b,a)\to(-\infty,\infty)$ with $a,b>0$, which is continuous, strictly increasing, and $\alpha(0)=0$. $\alpha$. Here $\dot{h}=\partial_{x}^{\top}h(x)\dot{x}$ is used to denote the time derivative of $h$ along the solution of (1), which results in:
@@
The link between the existence of a CBF and the forward invariance of the related safe set is established by the following key result, present in [Ames2017ControlSystems] and conveniently reported here applied to pH systems (1).
Theorem 1.
Let $h(x)$ be a CBF on $\mathcal{D}$ for (1). Any locally Lipschitz controller $u=k(x)$ such that $\{h,H\}_{J}-[h,H]_{R}+\partial_{x}^{\top}h(x)g(x)k(x)\geq-\alpha(h(x))$ provides forward invariance of $\mathcal{C}$. Additionally $\mathcal{C}$ is asymptotically stable on $\mathcal{D}$.
@@
The transformation of the desired control input $u_{\textrm{nom}}(x)$ in $u^{*}(x)$ by solving (6) is denoted as safety-critical control. A result that will be crucially used in this work is the following, presented in [Xu2015RobustnessControl] and presented here on pH systems (1) with the notation used in this paper.
Theorem 2.
Let $h(x)$ be a CBF on $\mathcal{D}$ for (1) and assume $\mathcal{U}=\mathbb{R}^{m}$. Define $\Psi(x):=\dot{h}(x,u_{\textup{nom}}(x))+\alpha(h(x))$ and assume $[h,h]_{gg^{\top}}\neq 0$ when $\Psi<0$. The Lipschitz continuous closed-form solution for (6) is given by $u^{*}(x)=u_{\textup{nom}}(x)+u_{\textup{safe}}(x)$, where
$u_{\textup{safe}}(x)=-\mathbb{I}_{\Psi<0}\frac{g^{\top}\partial_{x}h}{[h,h]_{gg^{\top}}}\Psi,$ (7)
where the indicator function $\mathbb{I}_{\Psi<0}$ returns $1$ if $\Psi<0$ ($u_{\textrm{safe}}\neq 0$) and 00 otherwise ($u_{\textrm{safe}}=0$).
Consistently with the arguments reported in Sec. II.A.4, in the remainder of the paper, (1) will represent a (possibly controlled) physical system, encoding desired closed-loop properties. This perspective, which is custom in energy-based frameworks (e.g., in IDA-PBC designs [ORTEGA2002585]), introduces a new perspective in the CBF framework, i.e., the desired closed-loop behavior is achieved with $u_{\textrm{nom}}=0$, which will be assumed throughout the paper.
@@ III Putting energy back in control through CBFs
In this section, we discuss how CBF-induced safety-critical control influences the power balance of the controlled physical system. The following result presents a closed-form solution for the power injected by the safety-critical controller (7).
@@
Under the assumptions of Theorem 2, the power balance for the pH system (1), undergoing safety-critical control (6) induced by a CBF $h(x)$, is
$\dot{H}=-[H,H]_{R}-\mathbb{I}_{\Psi<0}\frac{[H,h]_{gg^{\top}}}{[h,h]_{gg^{\top}}}\Psi,$ (8)
where $\Psi=\{h,H\}_{J}-[h,H]_{R}+\alpha(h(x))$.
@@
The results follows from an explicit calculation of the power balance (2) using $u=u_{\textup{safe}}(x)$ given by (7). The expression of $\Psi$ follows from (5) with $u=u_{\textrm{nom}}=0$.
@@
$P_{h,\alpha}=-\mathbb{I}_{\Psi<0}\frac{[H,h]_{gg^{\top}}}{[h,h]_{gg^{\top}}}\Psi$ (9)
where the notation stresses the fact that this term depends both on the CBF $h$ and on the function $\alpha$, a parameter of the CBF algorithm. We will now explore different control strategies induced by the analytic inspection of this power term.
III-1 CBF-based damping injection
As reported in Sec. II-A4, when the Hamiltonian $H$ qualifies as a Lyapunov function, the minima of $H$ form a stable set. The following result gives a useful sufficient condition for preservation of local stability, as well as a constructive method for novel damping injection schemes.
@@ III-1 CBF-based damping injection
Suppose $H$ is a Lyapunov function for the pH system (1). Under the assumptions of Theorem 2, let a safety-critical control algorithm induced by a CBF $h(x)$ act on the system. If $\mathbb{I}_{\Psi<0}[H,h]_{gg^{\top}}\leq 0$ in a neighborhood of the equilibrium, the stability of the original system is preserved, and the closed-loop system converges to the largest invariant subset in $\{x\in\mathcal{D}|P_{h,\alpha}-[H,H]_{R}=0\}$.
@@
The power balance for the safety-critically controlled system is given by $\dot{H}=-[H,H]_{R}+P_{h,\alpha}$. Using $H$ as a candidate Lyapunov function, a sufficient condition for local stability preservation of the original system is that in a neighborhood of the equilibrium $P_{h,\alpha}\leq 0$, which is equivalent to $\mathbb{I}_{\Psi<0}[H,h]_{gg^{\top}}\leq 0$. Due to time-invariance of closed-loop system, LaSalle’s invariance principle can be invoked to conclude that the system will converge to the largest invariant subset in $\{x\in\mathcal{D}|P_{h,\alpha}-[H,H]_{R}=0\}$. ∎
@@
The condition $P_{h,\alpha}\leq 0$ forces the safety-critical controller to “act as a damper", aiding dissipative dynamics in (1) in the stabilization task. This control strategy is normally referred to as damping injection in the pH framework. The closed-form expression (9) can be used to design and interpret the non-trivial damping effects that specific CBFs have on the controlled system, similarly to the analysis in [10136379], where passivity, rather than stability, has been addressed.
III-2 CBF-based energy-aware designs
The power term induced by CBFs (9) induces a constructive way to “put energy back in control", in the sense envisioned in energy-based methods (see e.g., [Ortega2001PuttingControl] and [Stramigioli2015Energy-AwareRobotics]), going beyond stabilization purposes and possibly injecting positive power in the system if the task requires it. Since energy is a function of the state of the system, CBFs offer a constructive tool for designing tasks that involve regulating the closed-loop energy, such as those required for achieving periodic locomotion, which cannot be achieved through purely passive designs. The following motivating example shows the concept, and highlights the role of the $\mathcal{K}$ function $\alpha$ in the algorithm.
Example 1 (Lower/Upper bound on total energy).